OpenStudy (anonymous):

4(a+2)=14-2(3-2a) A. all real numbers B. –1 C. –2 D. no solution

OpenStudy (joshawad):

first we distribute 4A + = 14- 6 - 4A

OpenStudy (joshawad):

4A + 8= 14- 6 - 4A

OpenStudy (cookiemonster2412):

all real numbers

OpenStudy (anonymous):

okay thank you! :))

OpenStudy (joshawad):

@cookiemonster2412 the point is to show them how to do the answer...